OpenStudy (anonymous):

A container has the dimensions of 10 cm x 50 mm x 0.2 m. It is filled completely with 5.0 kg of a powdery ore. What is the density of the ore in g/cm^3? (D = m/V) A. 0.05 g/cm^3 B. 5.0 g/cm^3 C. 50 g/cm^3 D. 500 g/cm^3 ***idk how to solve.. please explain?? thank youu!!

jimthompson5910 (jim_thompson5910):

first convert 5 kg to grams

OpenStudy (anonymous):

i get 5,000 grams??

jimthompson5910 (jim_thompson5910):

good, 1000 g = 1 kg, so 5000 g = 5 kg

OpenStudy (anonymous):

:) what do i do next?? and that is the mass right?? =D

jimthompson5910 (jim_thompson5910):

now convert each dimension to cm?

jimthompson5910 (jim_thompson5910):

yes we just converted the mass

jimthompson5910 (jim_thompson5910):

oh didn't mean to state that as a question lol

OpenStudy (anonymous):

(10cm)(5cm)(20cm) ??? did i do that right??

OpenStudy (anonymous):

oh haha its all good :) lol

jimthompson5910 (jim_thompson5910):

good, now multiply that out

OpenStudy (anonymous):

1000 cm??

jimthompson5910 (jim_thompson5910):

cubic cm or cm^3

jimthompson5910 (jim_thompson5910):

but yes, 1000 cm^3

OpenStudy (anonymous):

ohhh cm^3 then.... so now i plug into the equation?? d=m/v ??? so D=5000/1000 ???

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

so i get 5.0 g/cm ^3 ??? so my answer is B ?? :O

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

yay!! thank you :)

jimthompson5910 (jim_thompson5910):

yw